OpenStudy (anonymous):

Hw 12b, states I should 'break up the trajectory into two segments with the same time duration'. Does that mean I find the greatest distance traveled, find the max velocity based on that distance, split that duration of travel into two even parts and create an initial and final set of inverse kinematic vector equations for both the initial alpha and the final alpha? Meaning a total of four sets of three alpha vector equations, two sets for the initial alpha(0) and two sets for the final alpha(5)?

OpenStudy (anonymous):

You know the total time is 5 seconds for the trajectory. Two time segmets of equal duration would be 2.5 seconds each. You should find the end effector configuration at T=0, 2.5, and 5. Then figure out the joint angles at each time and use that to spline the positions together.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

because of 1a, I have the end configuration of the initial and the final configurations (T=0 and T=5). So I need to find the end configuration at T = 2.5

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Once I have all the end effector configurations, I will want to find the inverse kinematics so I will know all the joint angles at the three positions.

OpenStudy (anonymous):

Yeah but again you have the joint angles at T=0 and t=5 from the first part.

OpenStudy (anonymous):

Since I am to find the straight line trajectory and that means use the equation given.

OpenStudy (anonymous):

That will give me the midpoint end effector configuration.

OpenStudy (anonymous):

or what ever other time (t) I should want

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Finding the inverse kinematics of (x y theta)T sounded easier than I thought

OpenStudy (anonymous):

For the three link planar manipulator you solved that problem in the homework 2 weeks ago. You should have formulas for each joint angle.

OpenStudy (anonymous):

Yes, yes I did. I know x and y so I can find r. I know the lengths so that gives me the beta and gamma and delta. That gives me alpha1 and 2, and those give me alpha3. What is meant by the warning 'be careful of multiple solutions'. I know I will have multiple solutions for alpha1 and 2 but why the warning?

OpenStudy (anonymous):

r is actually a function of xw and yw dont forget that. I think the warning just means you should pick the solution that makes the most sense. In this case, I would imagine the alphas at the waypoint should be between those at T=0 and T=5. It's conceivable that they wouldn't be for one of the solutions.

OpenStudy (anonymous):

I think I did forget that about x and y because I am trying to figure out why I got a complex delta.

OpenStudy (anonymous):

The solutions are posted on t-square from the previous homeworks so you can get your formulas there.

OpenStudy (anonymous):

Gotta go for now.

OpenStudy (anonymous):

Ok thanks, later

OpenStudy (anonymous):

for 1b, I found the midpoint angles, formed the spline trajectory between initial and midpoint ( 3 functions) and the spline trajectory between midpoint and final (3 more functions). Is that all 1b is asking for? What does it mean by 'spline the desirded joint angles together'?

OpenStudy (anonymous):

This is essentially a spline with 1 waypoint(the midpoint). You should use what I talked about last thursday to spline all the angles together.

OpenStudy (anonymous):

I see the material we went over but I do not see how it it applies here.

OpenStudy (anonymous):

I am reading the notes but I do not follow how this applies.

OpenStudy (anonymous):

You have alphaf, alphai, alphav, tf, ti, and tv. You can make a splined trajectory with 1 waypoint.

OpenStudy (anonymous):

Will I have a p(t) and a q(t) for each set of alphas?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

how do you find the inverse kinematics of \[\xi \]

OpenStudy (anonymous):

In 1c, I am to do the same thing as in 1b but with \[\xi \] but how do I find the inverse of \[\xi \]

OpenStudy (anonymous):

Problem c is the same as problem b, except that instead of a linear end effector trajectory, the end effector trajectory is now defined by xi.

OpenStudy (anonymous):

this is SE(2), \[\omega \] is the sum of the main diagonal but I do not know g so I do not know \[\omega \] so I do not see how I can solve. This is what I have \[g _{e}\left( t \right)=g _{e}\left( 0 \right)e ^{\xi t}\] \[g _{e}^{-1}(0)g _{e}(t)=g _{e}^{-1}(0)g _{e}(0)e ^{\xi t}\] \[e ^{\xi t}=g _{e}^{-1}(0)g _{e}(t)\]

OpenStudy (anonymous):

You have ge(5), ge(0), and t. Isn't that all you need to solve for xi?

OpenStudy (anonymous):

I do know ge(0) and ge(5) and t=2.5 at the midpoint.

OpenStudy (anonymous):

if I made \[e ^{2.5 \xi }=g _{e}^{-1}(0)g _{e}(2.5)\] what would \[g _{e}(2.5) \] be

OpenStudy (anonymous):

Ok in part b you had a formula for the trajectory. You plugged in t=2.5 to get the ge at 2. correct?

OpenStudy (anonymous):

Am I not thinking about this in the right way?

OpenStudy (anonymous):

No you are not. Do you agree with my previous statement?

OpenStudy (anonymous):

Yes. I got \[g _{e}(2.5)\] by using the vector form of the problem. In part c I am instructed 'Instead of vectorizing' which tells me not to use the ge(2.5) from part b.

OpenStudy (anonymous):

Yes I just want you to agree that you got ge(2.5) by plugging in t=2.5 into a formula for the trajectory of ge.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok. In part c you also need a formula for the trajectory of the end effector. Once you get this formula, you can just plug in t = 2.5 to get ge(2.5). The only difference is this time there is a different formula for the trajectory of ge.

OpenStudy (anonymous):

\[g _{e}(t)=g _{e}(0)e ^{\xi t}\]

OpenStudy (anonymous):

Yes that's correct. What's the only unknown in that formula?

OpenStudy (anonymous):

\[\xi \]

OpenStudy (anonymous):

Ok. You should know how to get xi. It's been done in homeworks and you had to do it on the test.

OpenStudy (anonymous):

Yes, for SE(3). This is SE(2) and I am unsure as how to find \[\omega \]

OpenStudy (anonymous):

I have \[\omega = \left( \frac{ 1 }{ \tau } \right) atan(R _{21},R _{11})\] for \[R _{21} and R _{11}\] what rotation matrix do I create?

OpenStudy (anonymous):

Check homework 8 problem 3 solutions. You've done this in SE2

OpenStudy (anonymous):

To figure out xi, you first need g, the transformation from gi to gf.

OpenStudy (anonymous):

My mistake, I see that I did an SE(2) to find \[\xi \] I see how to find \[\xi \] now.

OpenStudy (anonymous):

I solve \[g _{e}(2.5) = g _{e}(0)e ^{\xi *2.5}\] and find the joint angles for \[g _{e}(2.5)\] then do what I did in part b

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thank you for your help, goodnight

OpenStudy (anonymous):

Good night you're welcome.

OpenStudy (anonymous):

If \[\xi = \left(\begin{matrix}x \\ y \\ z\end{matrix}\right)\] would \[e ^{\xi t} = e ^{\left(\begin{matrix}x \\ y \\ z\end{matrix}\right) t}=\left(\begin{matrix}e ^{xt} \\ e ^{yt} \\ e ^{zt}\end{matrix}\right)\]

OpenStudy (anonymous):

No. You need to take the matrix exponential. This has also been done numerous times before, including on the test.

OpenStudy (anonymous):

\[g _{e}(midway) = e ^{\xi \frac{ \tau }{ 2}} = \left[\begin{matrix}R(\omega \frac{ \tau }{ 2 }) & (-\frac{ 1 }{ \omega })(unit matrix - R(\omega \frac{ \tau }{ 2 }) )Jv \\ 0 & 1\end{matrix}\right]\] \[\xi = \left(\begin{matrix}v \\ \omega\end{matrix}\right) \] has been found

OpenStudy (anonymous):

Seems reasonable except you need to multiply by ge(0) on the left.

OpenStudy (anonymous):

I got the above equation from the notes. I believe I see why I would multiply it by ge(0) though. It is because of the trajectory we are trying to follow given in the problem

OpenStudy (anonymous):

Yes as before \[g_f = g_ie^{\xi*\tau}\]

OpenStudy (anonymous):

Is it \[g = g _{f}g _{i}^{-1}\] or \[g = g _{i}^{-1}g _{f}\] I have it in my notes both ways. I have done homework both ways and on my test one way

OpenStudy (anonymous):

I say it is \[g = g _{i}^{-1}g _{f}\]

OpenStudy (anonymous):

That's fine.

OpenStudy (anonymous):

1d asks to plot. I am not sure how to plot this.

OpenStudy (anonymous):

You've plotted trajectories before. What is different about this situation?

OpenStudy (anonymous):

I am just not sure of how to handle the p(t) and the q(t).

OpenStudy (anonymous):

With the planarR3 from the last homework, I need alpha 1, 2 and 3 but I have p(t) and q(t).

OpenStudy (anonymous):

Well first of all you don'e need to plot the planarR3, you just need to plot the frame of the end effector over time. You should have p(t) and q(t) for each alpha so for any time t, you can figure out alpha1, alpha2, and alpha3.

OpenStudy (anonymous):

are the p(t) and q(t) the alphas?

OpenStudy (anonymous):

I plot my end effector path from 0 to 2.5 using the p(t) equation to give me alpha1, 2 and 3, and from 2.5 to 5 using mthe q(t) to give me alpha1, 2 and 3. I use planarR3 and plug in the alpha's. My angles at p(2.5) do not = my angles at q(2.5) so I have a large displacement in the middle of my graph. Am I thinking about this right?

OpenStudy (anonymous):

Is 1c suppose to supply alpha values that stay between the initial and final configurations when plotted in 1d?

OpenStudy (anonymous):

Ok let's just do one question at a time. p(0) should equal q(0). The spline from T=2.5 to T=5 is equal to q(t-tv). So it should go from 0 to 2.5 just like p(t) does.

OpenStudy (anonymous):

p(2.5) should equal q(0) is what I meant. apologies

OpenStudy (anonymous):

I do understand and agree that p(2.5) should = q(0). I am plotting q(t) from 2.5 to 5, I think this might be my problem but I need to check to see if my p(2.5) = q(0).

OpenStudy (anonymous):

Yeah you should plot q(t-tv) from 2.5 to 5. That would be the issue.

OpenStudy (anonymous):

My p(2.5) does not = q(0). When I print p(1) I get three numbers but I thought I would only get one

OpenStudy (anonymous):

A problem the same as this, with three joints, was done in class. I wrote down a 8x3 matrix that was multiplied by a 3x1 alpha matrix the result being a 8x1 coeff matrix. I double check that I copied this down correctly and put it into matlab right. I use these coeff to complete the p(t) and q(t) functions that make the alphas. Is that much right in steps.

OpenStudy (anonymous):

Yes all that makes sense.

OpenStudy (anonymous):

Once you have the coefficients all you have to do is plot them.

OpenStudy (anonymous):

plot the trajectories that is.

OpenStudy (anonymous):

so plot p(t) for 0<t<2.5 and plot q(t-tv) for 2.5<t<5

OpenStudy (anonymous):

Ok I have to go for now. I'll check back later.

OpenStudy (anonymous):

do you think you will be back tonight?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

% p(t) for 0<t<2.5 alpha1p = -0.524 + 0*(2.5) + 0.0522*(2.5)^2 + -0.078*(2.5)^3 % q(t-tv) for 2.5<t<5 alpha1q = -0.284 + 0.1571*(2.5 - 2.5) +... 0.0107*(2.5 - 2.5)^2 + -0.0112*(2.5 - 2.5)^3 This is matlab code. The coeff are what I calculated in matlab. This is just alpha1 for p(2.5) and q(2.5 - 2.5) = q(t - tv) = q(0), they do not equal. None of the alphas at this point equals.

OpenStudy (anonymous):

for part b and part c should I have found g wrist before we found alpha1, 2 and 3 since each joint has only one axis of rotation?

OpenStudy (anonymous):

For parts b and c, you should find inverse kinematics exactly the same way you did 2 homeworks ago. First find xw,yw...calculcate alpha1 and alpha2, and then calculate alpha3. It's exactly the same problem as 2 homeworks ago.

OpenStudy (anonymous):

What does it mean when I get beta and delta as complex numbers?

OpenStudy (anonymous):

and my p(2.5) does not = q(0)

OpenStudy (anonymous):

For any of the alphas?

OpenStudy (anonymous):

No alphas equal between p(2.5) and q(0)

OpenStudy (anonymous):

I looked at your answers. I think all you had to do was subtract or add 2*pi to your angles from q.